Please help me.
Solve by completing the square.
2x^2+10x+11=0
I get-5+-sqrt23/2
not sure how I got there. Is it right? How should I work it out.

Answer by stanbon(75887)

Solve by completing the square.
2x^2+10x+11=0
x^2+5x+? = -11/2 +?
x^2+5x+(5/2)^2 = (-11/2)+(25/4)
(x+(5/2))^2 = 3/4
Take the square root of both sides to get:
x+(5/2)=+- (1/2)sqrt3
x= -(5/2)+-(1/2)sqrt3
Cheers,
Stan H.